1) Suppose that on a small island off the coast of Scotland, 32 percent of the population has blue eyes, which means that these individuals must be homozygous for the blue eye color gene (bb). The only other eye color found on the island is brown, and individuals that are homozygous for the brown eye color gene (BB) or heterozygous (Bb) will have brown eyes because brown is the dominant gene. Assume this population is in Hardy-Weinberg equilibrium. If 100 babies are born next year, how many of these would you expect to have brown eyes and be heterozygous?

a) 58

b) 49

c) 29

d) 43

3) Beak Length in Finches

In a certain population of finches, a single gene controls beak length. There are two possible alleles: S, which leads to a short beak, and L, which leads to a long beak. Individuals with the genotype SS will have a short beak, individuals with the genotype LL will have a long beak, and individuals with the genotype SL will have a medium-length beak. Scientists measure that in this population of 1,128 total birds, 321 birds have short beaks, 562 have medium-length beaks, and 245 have long beaks.

Ten years later, the measurements are repeated. This time, there were 1,346 birds in the population. 452 have short beaks, 699 have medium-length beaks, and 195 have long beaks.

Based on the original data collected from the population of 1,128 birds, what fraction of the population would be expected to have the SS genotype if the population was in Hardy-Weinberg equilibrium?

a) 0.466

b) 0.534

c) 0.285

d) 0.217

4) Beak Length in Finches

In a certain population of finches, a single gene controls beak length. There are two possible alleles: S, which leads to a short beak, and L, which leads to a long beak. Individuals with the genotype SS will have a short beak, individuals with the genotype LL will have a long beak, and individuals with the genotype SL will have a medium-length beak. Scientists measure that in this population of 1,128 total birds, 321 birds have short beaks, 562 have medium-length beaks, and 245 have long beaks.

Ten years later, the measurements are repeated. This time, there were 1,346 birds in the population. 452 have short beaks, 699 have medium-length beaks, and 195 have long beaks.

Based on the data collected 10 years later, is this population of finches currently evolving? How do you know?

a) yes because more individuals in the population exhibit the SS genotype than would be expected if the population was in Hardy-Weinberg equilibrium

b) no because the number of individuals that exhibit the SL genotype matches what would be expected if the population was in Hardy-Weinberg equilibrium

c) no because fewer individuals in the population exhibit the SS genotype than would be expected if the population was in Hardy-Weinberg equilibrium

d) yes because more individuals in the population exhibit the LL genotype than would be expected if the population was in Hardy-Weinberg equilibrium.
Answer #1

1)

Given

Blue eyes (bb)=32%=32/100=.32

BB, Bb = Brown Eyes

Out of 100, 32 will have blue eyes

so bb=32/100=.32

q2=.32

q=.phpKPMRgR.png=0.566

p+q=1

p+0.566=1

p=1-0.566=0.434

Hardy Weinberg equation=p2+q2+2pq=1

Heterozygous with blue eyes=Bb=2pq

=2×0.434×0.566

=0.4913

Percentage of babies=100×0.4913

=49%